2) If an electron in copper near the lowest possible temperature has a kinetic energy of 6.6*10^-19, what is the speed of the electron ?
1
Expert's answer
2019-08-07T10:01:08-0400

The kinetic energy

"KE=\frac{mv^2}{2}"

So, the speed of the electron

"v=\sqrt{\frac{2KE}{m}}=\sqrt{\frac{2\times 6.6\times 10^{-19}\:\rm{J}}{9.1\times 10^{-31}\:\rm{kg}}}"

"=1.2\times 10^6\:\rm{m/s}"